OpenStudy (anonymous):

What is the value of the expression 4/2^-2 ? A. ½ B. 1 C. 8 D. 16

OpenStudy (math&ing001):

\[\frac{ 4 }{ 2^{-2} } = 4 * 2^{2} = 16\]

OpenStudy (haseeb96):

D is the correct answer

OpenStudy (anonymous):

D IS THE CORRECT ANSWER

OpenStudy (haseeb96):

When the power on any base is negative at the denominator then when we put on the numenatore then its -ve sign change into +be sign.and it will become as shown by the above user .